Circular No. 6438

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996al IN NGC 7689 S. Benetti, European Southern Observatory (ESO); and R. Neuhauser,

Max-Planck-Institut, Garching, report: “Inspection of a preliminarily-re-duced CCD spectrum (range 440–705 nm, resolution 0.35 nm) obtained on July 23.35 UT with the ESO 1.52-m telescope (+ Boller & Chivens spectrograph) at La Silla confirms this object as a supernova of type II. Strong emission lines of Hβ, He I (587.6 nm), and Hα, with complex profiles, are superimposed on a fairly blue continuum. Hβ consists of a broad emission (FWHM ∼ 15 300 km/s), on which is superimposed a narrow P-Cyg line with an expansion velocity of 784 km/s (if a recession velocity of 1981 km/s is assumed for NGC 7689). He I at 587.6 nm consists of a very broad, flat-topped emission (FWHM ∼ 22 900 km/s), on which is superimposed a narrow, unresolved emission line (centered at 591.57 nm). Hα consists of a very broad emission (FWHM ∼ 20 400 km/s), a broad emission (FWHM ∼ 9100 km/s), and a narrow P-Cyg component (expansion velocity 850 km/s) — dominated by a strong emission peaking at 660.97 nm (FWHM = 0.95 nm) — that is superimposed on the broad emission. The Fe II 492.4-, 501.8-, and 516.9-nm lines are also present as narrow absorptions (FWHM ∼ 0.8 nm), with the minima blueshifted by 0.4 nm with respect to the rest-frame positions. The characteristics of this spectrum are those of a type-II supernova caught soon after explosion; the spectrum also shows signs of circumstellar material probably lost by the supernova progenitor before exploding. The spectrum of SN 1996al is expected to evolve quite rapidly, and monitoring of this very interesting supernova is badly needed at all wavelengths.”

COMET C/1996 B2 (HYAKUTAKE) Total visual magnitude estimates (cf. IAUC 6420): June 19.82 UT, 5.8

(D. Seargent, Cowra, N.S.W., naked eye); 24.77, 6.3 (Seargent, The Entrance, N.S.W., 25×100 binoculars); July 2.35, 6.4 (P. Camilleri, Cobram, Victoria, Australia, 20×80 binoculars); 7.39, 6.9 (Camilleri); 19.29, 7.6 (A. Pearce, Wanaka, New Zealand, 20×80 binoculars); 20.30, 7.6 (Pearce).

SUPERNOVA 1996ak IN NGC 5021 Corrigendum. On IAUC 6434, line 2, for 2.12-m read 0.20-m

1996 July 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green